To estimate animal populations, biologists count the total number of animals in a small section of a habitat. The total population of animals is directly proportional to the size of the habitat (in acres)polled.
There are 3 parts to the question and I'm struggling with setting them up.
a. Write an equation using only one variable that could be used to solve for the constant of variation k.
b. A biologist counted the number of white tail deer in a 100 acre parcel of land in a nature preserve and found 12 deer. Find the constant of variation k.
c. If the entire nature preserve is 2,500 acres, then what is the total white tail deer population in the preserve? Describe how you arrived at your answer.
I would really appreciate someone explaining how to work these 3 parts of the problem. I have several other similar ones and don't think I'll understand those until I understand this one. Thanks for your help!

Answer by stanbon(75887)

To estimate animal populations, biologists count the total number of animals in a small section of a habitat. The total population of animals is directly proportional to the size of the habitat (in acres)polled.
-------
population = k*habitat size
p = k*a
---------------------
There are 3 parts to the question and I'm struggling with setting them up.
a. Write an equation using only one variable that could be used to solve for the constant of variation k.
p = k*a
-----------------------------
b. A biologist counted the number of white tail deer in a 100 acre parcel of land in a nature preserve and found 12 deer. Find the constant of variation k.
-------
12 = k*100
k = 0.12
---------------
c. If the entire nature preserve is 2,500 acres, then what is the total white tail deer population in the preserve? Describe how you arrived at your answer.
------
p = 0.12*a
p = 0.12*2500 = 300
========================
Cheers,
Stan H.